# Altogether: Image Captioning via Re-aligning Alt-text

### v1

@@ 4.2 Captioner Architecture
We choose the pretrained MetaCLIP ViT-H/14 Xu et al. (2024) as the image encoder, which outputs a single embedding with 1024 dimensions. The image embedding is then transformed into 40 visual tokens via the mapping network to serve as the image prompt for the text decoder. We freeze the image encoder during the training phase and only train the mapping network.
@@ 5.3 Text-to-image (T2I) Generation
MetaCLIP 72.4 80.5 94.2 98.0 86.4 83.4 74.1 90.0 50.2 72.4 95.4 95.6 85.1 72.7 55.2 99.4 66.3 74.6 62.5 38.2 37.2 65.8 82.2 64.1 30.1 59.3 69.2
Altogether(3) ($p{=}0.15$) 73.2 82.1 95.0 97.8 87.1 88.6 74.6 93.1 63.2 73.0 95.9 95.9 86.8 86.1 54.6 99.5 70.3 76.0 57.9 28.1 43.3 50.1 85.4 65.4 32.5 58.3 62.5
Table 6: Results on 26 CLIP zero-shot classification tasks. First section: Training with pure ($p{=}1.0$) synthetic captions from our captioners that were trained after different rounds of annotations. Second section: Mixing in alt-text during training (ratio of $p{=}0.15$). Third section: Comparison of a large ViT-H/14 model trained on our synthetic captions with mixed alt-text outperforms MetaCLIP Xu et al. (2024) (72.4 vs. 73.2 average accuracy).
@@
MetaCLIP 60.4 85.0 57.5 66.1 64.6
Altogether(3) ($p{=}0.15$) 65.7 87.6 60.7 67.3 66.2
@@ Results.
The next section of Table 6 shows that training with only alt-text performs better than using only synthetic captions above. We believe this is because the captioner is likely not large enough to carry all the alt-text information into the synthetic caption. We then mix alt-text and synthetic captions (ablation in Appendix §D) for training CLIP. With a ratio of $p{=}0.15$ synthetic captions, we see a +1.1% improvement over 26 classification tasks (Table 6), showing how re-align can provide complementary information for CLIP training. Finally we train a large ViT-H/14 model with mixed Altogether captions and observe 73.2% average accuracy compared to the 72.4% with the same model in MetaCLIP Xu et al. (2024).
Finally, we evaluate on zero-shot text-to-image retrieval tasks from DataComp (Gadre et al., 2023). Results are in Table 7. Mixing alt-text with synthetic captions leads to +3% for retrieval on ViT-B and even larger gains over MetaCLIP ViT-H/14.
@@ Appendix A Annotation Guidelines
Frozen Encoder MetaCLIP Xu et al. (2024)
@@
Table 12: Hyperparameters of CLIP training.

### v2

@@ 4.2 Captioner Architecture
We choose the pretrained Meta CLIP 1.0 ViT-H/14 Xu et al. (2024) as the image encoder, which outputs a single embedding with 1024 dimensions. The image embedding is then transformed into 40 visual tokens via the mapping network to serve as the image prompt for the text decoder. We freeze the image encoder during the training phase and only train the mapping network.
@@ 5.3 Text-to-image (T2I) Generation
Meta CLIP 1.0 72.4 80.5 94.2 98.0 86.4 83.4 74.1 90.0 50.2 72.4 95.4 95.6 85.1 72.7 55.2 99.4 66.3 74.6 62.5 38.2 37.2 65.8 82.2 64.1 30.1 59.3 69.2
Meta CLIP 1.2 (Altogether(3), $p{=}0.15$) 73.2 82.1 95.0 97.8 87.1 88.6 74.6 93.1 63.2 73.0 95.9 95.9 86.8 86.1 54.6 99.5 70.3 76.0 57.9 28.1 43.3 50.1 85.4 65.4 32.5 58.3 62.5
Table 6: Results on 26 CLIP zero-shot classification tasks. First section: Training with pure ($p{=}1.0$) synthetic captions from our captioners that were trained after different rounds of annotations. Second section: Mixing in alt-text during training (ratio of $p{=}0.15$). Third section: Comparison of a large ViT-H/14 Meta CLIP 1.2 model trained on our synthetic captions with mixed alt-text outperforms Meta CLIP 1.0 Xu et al. (2024) (72.4 vs. 73.2 average accuracy).
@@
Meta CLIP 1.0 60.4 85.0 57.5 66.1 64.6
Meta CLIP 1.2 (Altogether(3), $p{=}0.15$) 65.7 87.6 60.7 67.3 66.2
@@ Results.
The next section of Table 6 shows that training with only alt-text performs better than using only synthetic captions above. We believe this is because the captioner is likely not large enough to carry all the alt-text information into the synthetic caption. We then mix alt-text and synthetic captions (ablation in Appendix §D) for training CLIP. With a ratio of $p{=}0.15$ synthetic captions, we see a +1.1% improvement over 26 classification tasks (Table 6), showing how re-align can provide complementary information for CLIP training. Finally we train a large Meta CLIP 1.2 ViT-H/14 model with mixed Altogether captions and observe 73.2% average accuracy compared to the 72.4% with the same model in Meta CLIP 1.0 Xu et al. (2024).
Finally, we evaluate on zero-shot text-to-image retrieval tasks from DataComp (Gadre et al., 2023). Results are in Table 7. Mixing alt-text with synthetic captions leads to +3% for retrieval on ViT-B and even larger gains over Meta CLIP 1.0 ViT-H/14.
@@ Appendix A Annotation Guidelines
Frozen Encoder Meta CLIP 1.0  Xu et al. (2024)
@@
Table 12: Hyperparameters of Meta CLIP 1.2 training.